OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y − 3 = x2 + 4x? (−3, 0) and (−1, 0) (3, 0) and (1, 0) (−3, 0) and (1, 0) (3, 0) and (−1, 0)

OpenStudy (anonymous):

y = x^2 + 4x + 3 0 = x^2 + 4x + 3 0 = (x + 3)(x + 1) x = -3, -1

OpenStudy (anonymous):

So, it is the first one

OpenStudy (anonymous):

Thanks.